OpenStudy (anonymous):

x^2-9x, x^2+9x find LCM

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

LCM : x(x-9)(x+9)

OpenStudy (anonymous):

thats the final answer?

OpenStudy (anonymous):

taking commons out, \[x^{2}-9x=x(x-9)\] and \[x^{2}+9x=x(x+9)\] now LCM=(common terms in both)*(remaining terms)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

??

geerky42 (geerky42):

LCM=(common terms in both)*(remaining terms) Both have x. and the remaining is (x-9) and (x+9) So \(\text{LCM} = \boxed{x(x-9)(x+9)}\)